TEMPERATURE MEASUREMENTS IN VOLUMETRIC MICROWAVE DISCHARGES ^{a}

\underline{J.M. LEWIS}, I. ADAMOVICH, W.R. LEMPERT, Department of Mechanical Engineering, The Ohio State University, Columbus, OH 43210.

The microwave discharge has been proposed as an excitation source to create and maintain stable, large-volume low to moderate temperature plasmas at near atmospheric pressure. We study a microwave discharge produced from a variety of combinations of flowing Ar, CO and N_{2}. We concentrate on rotational temperature measurements based on spontaneous emission spectra from radicals such as C_{2} and CN. Spectroscopic measurements are complimented by stability and uniformity measurements obtained by short exposure time-resolved images from an intensified CCD camera. The feasibility of using a microwave source for creating a large-volume plasma based on these temperature measurements, energy requirements, and basic plasma characterization will be reported.

aResearch is primarily funded by AFOSR and DDRE under the Air Plasma Rampart MURI